# Sensing-Enhanced Channel Estimation for Near-Field XL-MIMO Systems

### v1

@@ VI-D Storage and Baseband Sample Analysis
Although the spherical wave dictionary achieves faster convergence during the initial iterations, the convergence performance is not as good as the other dictionaries due to the mutually correlated codewords. Besides, the two DoFs in both distance and angle of the spherical wave dictionary dramatically add to the dictionary size as the resolution requirement increases. Therefore, the spherical wave dictionary requires the largest number of baseband samples and codewords to reach the target NMSE. On the contrary, thanks to the mutual orthogonality among codewords, the DFT dictionary can satisfy more stringent NMSE requirements with smaller sizes than the spherical ones. Yet, its mismatch with the near-field channel model still leads to a bulkier dictionary compared to the proposed DPSS-based one.
In contrast, thanks to the prior localization information integrated in the proposed dictionary, it is able to reach the NMSE targets with a drastically smaller number of baseband samples. For example, to achieve a $-20\,$dB NMSE, the required baseband samples show a $51\%$ and $77\%$ reduction compared to the DFT and spherical dictionaries, respectively. Furthermore, compared to the DFT and spherical wave dictionaries, the proposed DPSS-based dictionary does not need to sacrifice the NMSE performance for a lower storage, which is exemplified by an $88\%$ reduction in dictionary size for a given $-20\,$dB target NMSE.
@@ Appendix A Proof of Proposition 2
If the phase shift of reference wave is drawn from a uniform distribution, i.e., $\xi_{n}\sim{\mathcal{U}}[0,\pi]$, the interference term will vanish to 00 asymptotically, as $\mathbb{E}[\cos(2\xi_{n})]=0$ and $\mathbb{E}[\sin(2\xi_{n})]=0$ hold for $N_{\rm BS}\rightarrow\infty$. ∎

### v3

@@ VI-D Storage and Baseband Sample Analysis
Although the spherical wave dictionary achieves faster convergence during the initial iterations, the convergence performance is not as good as the other dictionaries due to the insufficient distance sampling interval and mutual correlation among codewords. Besides, the two DoFs in both distance and angle of the spherical wave dictionary dramatically add to the dictionary size as the resolution requirement increases. Therefore, the spherical wave dictionary requires the largest number of baseband samples and codewords to reach the target NMSE. On the contrary, thanks to the mutual orthogonality among codewords, the DFT dictionary can satisfy more stringent NMSE requirements with smaller sizes than the spherical ones. Yet, its mismatch with the near-field channel model still leads to a bulkier dictionary compared to the proposed DPSS-based one.
In contrast, thanks to the prior localization information integrated in the proposed dictionary, it is able to reach the NMSE targets with a drastically smaller number of baseband samples. For example, to achieve a $-20\,$dB NMSE, the required baseband samples show a $51\%$ reduction compared to the DFT and spherical dictionaries, respectively. Furthermore, compared to the DFT and spherical wave dictionaries, the proposed DPSS-based dictionary does not need to sacrifice the NMSE performance for a lower storage, which is exemplified by a $66\%$ reduction in dictionary size for a given $-20\,$dB target NMSE.
@@ Appendix A Proof of Proposition 2
If the phase shift of reference wave is drawn from a uniform distribution, i.e., $\xi_{n}\sim{\mathcal{U}}[0,\pi]$, the interference term will vanish to 00 asymptotically, as $\mathbb{E}[\cos(2\xi_{n})]=0$ and $\mathbb{E}[\sin(2\xi_{n})]=0$ hold for $N_{\rm BS}\rightarrow\infty$.[37]